OpenStudy (brixzii):

If m varies directly as the square root of y, inversely as p^2, and directly as n, what happens to m when y is quadrupled, p is tripled, and n is multiplied by 5?

OpenStudy (anonymous):

\[m \alpha \frac{ \sqrt{y} }{ p^{2} }(n)\] ----- \[m \alpha \frac{ \sqrt{(4)y} }{ (3p)^{2} }(5n)\]

OpenStudy (anonymous):

\[m \alpha (\frac{ 10 }{ 9 })[\frac{ \sqrt{y} }{ p^{2} }(n)]\]